# SpotVLM: Cloud-edge Collaborative Real-time VLM based on Context Transfer

### v1

@@ 6 Conclusion
In this work, we propose a novel context transfer paradigm that enables LVLMs to provide real-time guidance to SVLMs through semantic and visual contexts. Building on this paradigm, we introduce SpotVLM, a collaborative framework that integrates two corresponding modules: Context Replacement Module and Visual Focus Module. Extensive experiments across diverse tasks, datasets, and model combinations demonstrate that SpotVLM outperforms existing collaboration methods. Moving forward, we envision integrating stronger foundation models and adaptive scheduling strategies into the collaborative framework to further optimize performance in real-world deployments.

### v2

@@ 5 Conclusion
In this work, we propose a novel context transfer paradigm that enables LVLMs to provide real-time guidance to SVLMs through semantic and visual contexts. Building on this paradigm, we introduce edgeVLM, a collaborative framework that integrates two corresponding modules: Context Replacement and Visual Focus. Extensive experiments across diverse tasks, datasets, and model combinations demonstrate that edgeVLM outperforms existing collaboration methods. Moving forward, we envision integrating stronger foundation models and adaptive scheduling strategies into the collaborative framework to further optimize performance in real-world deployments.